OpenStudy (ksaimouli):

A cheetah, the fastest of all land animals over a short distance, accelerates from rest to 26m/s. Assuming that the acceleration is constant, find the average speed of the cheetah

OpenStudy (anonymous):

If acceleration is constant, then velocity grows linearly and you can take average of initial and final velocities.

OpenStudy (ksaimouli):



OpenStudy (anonymous):

Yep. You can verify it using definition of average velocity = total distance ÷ total time. The total distance, x=0.5(a)(t)^2, where 'a' is the constant acceleration and 't' is the time. You can make up whatever number you want for 't' and find 'a' by (final velocity) ÷ (total time).

OpenStudy (amistre64):

i do not see a "total distance" in the question

OpenStudy (anonymous):

@amistre64 q.v. "You can make up whatever number you want for 't' and find 'a' by (final velocity) ÷ (total time)."

OpenStudy (amistre64):

im just not sure how we can come up with a definitive solution with the information given.

OpenStudy (anonymous):

e.g. Let t = 3s. a = 26m/s ÷ 3s ≈ 8.67m/s/s. x = (0.5)(8.67m/s/s)(3s)^2 = 39m. average velocity = 39m/3s = 13m/s.

OpenStudy (anonymous):

or average velocity = (final v - initial v)/2. [because the velocity increase is linear].

OpenStudy (amistre64):

ok, thats starting to make better sense to me now

OpenStudy (anonymous):

We have initial velocity, final velocity, and the fact that acceleration is constant. That is plenty of information.

Parth (parthkohli):

I hate aww snap.

OpenStudy (amistre64):

the average of the instantaneous velocities .... i was reading it askew :)

Parth (parthkohli):

OK, so we need \(t\) right now.\[v^2 = u^2 + 2as\]We could've calculated the acceleration through the formula we already know.

Parth (parthkohli):

Then we had one variable to solve for: s(or the distance).

Parth (parthkohli):

After solving for s, we count have found s/t which is the speed.

OpenStudy (anonymous):

How do you get s without a? You need t to find a. If you already have t and a, you can get s with s=.5at^2. (The fundamental kinematics formula).

OpenStudy (anonymous):

Or... you can just do (final - initial)/2.